# Adversarial Disentanglement by Backpropagation with Physics-Informed Variational Autoencoder

### v2

@@ 5 Synthetic case studies
Three synthetic case studies of different complexity, illustrated in Figure 1, are discussed in detail throughout this section. Each case addresses a different set of challenges. The beam case study demonstrates that the proposed approach preferentially utilizes the known physics, yielding an interpretable and parsimonious representation of the physical system. The oscillator case study highlights how the adversarial training can prevent the model from learning arbitrary components of the response in a data-driven manner, and investigates the impact of the GRL hyperparameter. Finally, the bridge case study demonstrates the feasibility of using the model for damage detection in a more complex synthetic case, and compares the performance to that of existing data-driven approaches. For the purposes of reproducibility, the code needed to replicate the examples is made available on GitHub [^3] https://github.com/JanKoune/DPI-VAE.
@@
For all the case studies presented in this section, $h_{\mathrm{c}}$ and $h_{\mathrm{y}}$ are taken as the identity function for simplicity. Furthermore we use $N_{\mathrm{train}}=1024$ and $N_{\mathrm{val}}=512$, and consider no other regularization except for the GRL, i.e. $\beta=\alpha_{\mathrm{x}}=\alpha_{\mathrm{c}}=\alpha_{\mathrm{y}}=1.0$. To enable the formulation of problems with bounded latent variables and to ensure a stable optimization procedure, the physics-based latent variables are internally transformed to an unbounded and normalized workspace. All physics-grounded latent variables are constrained to lie within ranges that ensure consistency with the underlying physics of each system. Additional details regarding the architecture, data, optimization, and plotting are provided in Appendix A. It is noted that the case studies are only meant as didactic examples, intended to elaborate the issues with combining physics-based and data-driven components in VAE, provide intuition about the interaction between these components, and demonstrate the behavior of the model. Therefore, emphasis is placed on clarity rather than realism.
@@ 5.1 Beam case study
The case study consists of a beam with fixed length $L=1.0$ m and a point load with magnitude $F=1.0$ N acting on an uncertain position $x_{\mathrm{F}}$ along the length of the beam. The material is linear elastic, with uncertain Young’s modulus $E$, Poisson ratio $\nu=0.3$l area moment of inertia $I=2\cdot 10^{-6}$ m4 and $A=2.4\cdot 10^{-3}$ m2. The rotational stiffness of the right-hand side support is temperature dependent, with the dependence modeled as an increase in the rotational stiffness of the support at lower temperatures. The relationship between the temperature and the support rotational stiffness is formulated as $\log k_{\mathrm{r}}=8-\frac{10}{1+e^{-T/2}}$. The beam is subject to variability in the vertical stiffness of the right-hand side support, e.g. due to damage or a deficiency of the support, simulated as a translational spring boundary condition with stiffness $k_{\mathrm{v}}$. This quantity can span several orders of magnitude, and therefore we parametrize the model using $\log k_{\mathrm{v}}$ instead. The beam is equipped with $d_{\mathrm{x}}=32$ sensors measuring the vertical displacement, equally spaced along the length as shown in Figure 1(a).
where $b=L-x_{\mathrm{F}}$, and the non-bold $x$ refers to the position along the beam. This partial model represents the beam in the undamaged condition at a reference temperature. The nominal model is directly incorporated in the physics-based branch of the VAE decoder. The Young’s modulus $E$ and the position of the point load $x_{\mathrm{F}}$ are considered as uncertain latent variables, such that $\bm{z}_{\mathrm{x}}=(E,x_{\mathrm{F}})$. It is assumed that the temperature is an observed domain variable such that $\bm{c}=(T)$, and that the vertical spring log-stiffness $\log k_{\mathrm{v}}$ is taken as a class variable representing damage in the structure, such that $\bm{y}=(\log k_{\mathrm{v}})$. Since the class variable $\bm{y}$ represents damage in the structure it will not be quantitatively measurable. In a realistic scenario, observations of the condition of the support on a qualitative scale (e.g. from 00 representing no damage to $5$ denoting a fully damaged support) might be available. In this example we simplistically consider $\bm{y}$ as the ground truth value of $\log k_{\mathrm{v}}$ with some added noise. The variable symbols, units, types, as well as the prior distributions over the physics-based latent variables and the ground truth distributions of the generative factors used to generate the training data are summarized in Table 1. We additionally provide a reference value which is used to produce the figures and further discussed below. To ensure physical consistency and to avoid numerical issues, the Young’s modulus is truncated below a small positive value, and the load position $x_{\mathrm{F}}$ is restricted to the range $(0,1)$.
@@
Following the procedure described in Section 5, the full model is used to produce synthetic data by first drawing samples of the input parameters from the ground truth distribution, and subsequently contaminating the resulting model predictions with zero-mean Gaussian white noise with standard deviation $\sigma_{\mathrm{x}}=0.02$ m. The dataset used to train the VAE is composed of $N_{\mathrm{train}}$ measurements of the beam displacement $\bm{x}=\{\bm{x}_{i}\}_{i=1}^{N_{\mathrm{train}}}$ where each element $\bm{x}_{i}$ is a vector of length $d_{\mathrm{x}}=32$. The domain and class observables $\bm{c}=\{\bm{c}_{i}\}_{i=1}^{N_{\mathrm{train}}}$ and $\bm{y}=\{\bm{y}_{i}\}_{i=1}^{N_{\mathrm{train}}}$ are obtained as the ground truth values used to generate the dataset, with the addition of i.i.d. samples of Gaussian white noise with standard deviations of $\sigma_{c}=\sigma_{y}=0.02$ respectively. The dimensionality of the domain and class latent variables is taken as $d_{z_{\mathrm{c}}}=d_{z_{\mathrm{y}}}=2$.
After training, the disentanglement between physics-grounded and data-driven components is qualitatively assessed by examining the latent space and samples of the reconstructed response of the beam. A particularly useful tool for assessing the learned representation is to traverse the latent space and the space of reconstructions of the VAE. This can be achieved by generating synthetic data while interpolating over a specified generative factor, and setting the remaining generative factors to a constant reference value, provided in Table 1. The VAE is then evaluated on the generated data, yielding samples from the latent space, realizations of the reconstructed input $\hat{\bm{x}}$, and the mean physics-based and data-driven components $\hat{\bm{x}}_{\mathrm{p}}$ and $\hat{\bm{x}}_{\mathrm{d}}$. Each generative factor is linearly interpolated within the $1^{\mathrm{st}}$ and $99^{\mathrm{th}}$ percentiles of the corresponding ground truth distribution. Additional details on the plotting are provided in Appendix A.
@@
The predicted physics-based $\hat{\bm{x}}_{\mathrm{p}}$ and data-driven $\hat{\bm{x}}_{\mathrm{d}}$ components, as well as the combined prediction $\hat{\bm{x}}$ are shown in Figure 5. It can be observed that the data-driven component of the reconstruction $\hat{\bm{x}}_{\mathrm{d}}$ (middle row) is invariant to changes in $E$ and $x_{F}$ contributing only a constant deformed shape to the total predicted response. On the other hand, the physics-based model captures variability in both the physics-based generative factors $\bm{s}_{\mathrm{x}}=(E,x_{\mathrm{F}})$, but also domain and class generative factors $\bm{s}_{\mathrm{c}}=(T)$ and $\bm{s}_{\mathrm{y}}=(\log k_{\mathrm{v}})$. In contrast to the behavior of the unconstrained VAE, presented in Section 2.4, here the model preferentially utilizes the known physics. A key aspect of the adversarial training is the degree to which it allows interaction between the physics-based and data-driven components of the prediction. In this case study, the additional displacement of the beam due to the reduced vertical stiffness of the right-hand side support will also depend on the load position $x_{\mathrm{F}}$. More positive values of $\lambda$ tend to prevent the model from capturing this interaction, whereas more negative values enable it but may result in the data-driven components overriding the known physics.
@@
Figure 5: Mean prediction and ±$\pm 2\sigma$ uncertainty bounds for the physics-based $\hat{\bm{x}}_{\mathrm{p}}$ and data-driven $\hat{\bm{x}}_{\mathrm{d}}$ components, and combined prediction $\hat{\bm{x}}$ while traversing the generative factors. The input response measurements are denoted as dots in the bottom row.
To further highlight the impact of the GRL, the latent space traversals of the unconstrained model and the model trained adversarially are compared in Figure 6. Without adversarial training, the domain latent variables $\bm{z}_{\mathrm{c}}$ encode the variability in the load position $x_{\mathrm{F}}$, providing the data-driven decoder components with the information needed to reconstruct this component of the measured response and resulting in the issue discussed in Section 2.4. In contrast, when $\lambda=1/256$, the adversarial training results in a posterior distribution over $\bm{z}_{\mathrm{c}}$ that is invariant to changes in $x_{\mathrm{F}}$. The results shown previously indicate that the latent bottleneck architecture and GRL regularization result in a sparse and parsimonious representation of the physical system, and can yield domain and class latent variables that are invariant to changes in the underlying physics. The influence of the GRL hyperparameter $\lambda$ is further investigated in the oscillator case study presented below.
@@ 5.2 Oscillator case study
This example demonstrates how the adversarial training prevents the model from compensating for all discrepancies between the physics-based model predictions and measurements. Suppose that a mass-spring-dashpot system undergoes damped harmonic motion, starting from an initial displaced position $x_{0}$, with no external excitation. It is assumed that each experiment is performed under varying temperature $T$, which is taken as the domain variable. The temperature affects the spring stiffness through the relationship $k(T)=k_{\mathrm{ref}}+\alpha_{\mathrm{T}}(T_{\mathrm{ref}}-T)$, with $T_{\mathrm{ref}}=20.0$ Co and $\alpha_{\mathrm{T}}=0.01$. The reference spring stiffness at $T_{\mathrm{ref}}=20.0$ Co is assumed known and equal to $k_{\mathrm{ref}}=1.0$ N/m. The mass $m$ is considered unknown and treated as a physics-grounded latent variable to be inferred from data. The viscous damping coefficient $c_{\mathrm{d}}$ is taken to define the class of the system. Finally, it is assumed that the observations are subject to an unknown confounding influence in the form of small random perturbations of the initial displacement $x_{0}$. The equation of motion describing the system can be written as:
Each triplet of observations is composed of a noisy displacement time series, and noisy measurements of the viscous damping coefficient $c_{\mathrm{d}}$ and temperature $T$, which are considered as class and domain variables respectively such that $\bm{c}=(T)$ and $\bm{y}=(c_{\mathrm{d}})$. Training and validation datasets are generated by drawing samples from the ground truth distribution and generating the oscillator displacement time-series using the equation of motion shown in Equation 5.2. The synthetic response measurements are subsequently contaminated with i.i.d. realizations of Gaussian white noise with standard deviation $\sigma_{\mathrm{x}}=0.01$ m. The standard deviation of the measurement uncertainty of the domain and class observables are taken as $\sigma_{\mathrm{c}}=0.01$ and $\sigma_{\mathrm{y}}=0.01$ respectively. A summary of the generative factors, the prior distribution and the ground truth distribution is provided in Table 2. Each of the measured time-series is a vector of $64$ measurements, equally spaced within a time interval $t\in[0,10]$ s.
@@
The domain and class latent space dimensions are taken as $d_{z_{\mathrm{c}}}=d_{z_{\mathrm{y}}}=4$. The proposed model with no adversarial training ($\lambda=-1$) is trained and evaluated on the synthetic example. The reconstructed response obtained from a traversal of the initial displacement $x_{0}$, shown in 7(a), highlights another issue that occurs when combining physics-based and data-driven components in VAE: Although the variability in $x_{0}$ can not be accounted for by the physics-based model, and there is no information in the domain or class variables regarding the value of $x_{0}$, the lack of regularization results in a model that is free to capture the components of the measured displacement stemming from the variability in the initial displacement $x_{0}$. Although in this case the effect is benign, in more complex physical systems it can result in the model learning unknown confounding influences in a non-interpretable black-box manner. When the GRL regularization is utilized (7(b)), the data-driven encoder is unable to capture the variability in $x_{0}$, depriving the data-driven decoder from the information needed to reconstruct this component of the input measurements.
(a) Mean and ±$\pm 2\sigma$ uncertainty bounds of the reconstructed response without adversarial training ($\lambda=-1$).
(b) Mean and ±$\pm 2\sigma$ uncertainty bounds of the reconstructed response with adversarial training ($\lambda=1/128$).

### v3

@@ 5 Synthetic case studies
Three synthetic case studies of different complexity, illustrated in Figure 1, are discussed in detail throughout this section. The case study objectives, the definition of the physics-based models, the procedure used to generate the synthetic data, and details of the model implementation and visualization are provided below. For the purposes of reproducibility, the code needed to replicate the examples is made available on GitHub [^3] https://github.com/JanKoune/DPI-VAE. Additional information regarding the architecture, variable transformations, data, optimization, and visualization is provided in Appendix A.
Case study objectives
Each case study addresses a different set of challenges. The beam case study demonstrates that the proposed approach preferentially utilizes the known physics, yielding an interpretable and parsimonious representation of the physical system. The oscillator case study highlights how the adversarial training can prevent the model from learning arbitrary components of the response in a data-driven manner, and investigates the impact of the GRL hyperparameter. Finally, the bridge case study demonstrates the feasibility of using the model for damage detection in a more complex synthetic case, and compares the performance to that of existing data-driven approaches. It is noted that the case studies are only meant as didactic examples, intended to elaborate the issues with combining physics-based and data-driven components in VAE, provide intuition about the interaction between these components, and demonstrate the behavior of the model. Therefore, emphasis is placed on clarity rather than realism.
Physics-based models
@@
Implementation details
For all the case studies presented in this section, $h_{\mathrm{c}}$ and $h_{\mathrm{y}}$ are taken as the identity function for simplicity. Furthermore we use $N_{\mathrm{train}}=1024$ and $N_{\mathrm{val}}=512$, and consider no other regularization except for the GRL, i.e. $\beta=\alpha_{\mathrm{x}}=\alpha_{\mathrm{c}}=\alpha_{\mathrm{y}}=1.0$. Unless stated otherwise, the number of the domain and class latent variables are taken to be twice the number of domain and class generative factors [^4] The intention behind this choice is to avoid biasing the model towards a disentangled representation by matching the number of latent variables to the ground truth generative factors, ensuring that any disentanglement in the learned representation is not a consequence of limited latent space capacity. To enable the formulation of problems with bounded latent variables and to ensure a stable optimization procedure, the physics-grounded latent variables are obtained through a sequence of invertible transformations applied to the encoder output, mapping samples from an unbounded base latent space to the target latent space. All physics-grounded latent variables are constrained to lie within ranges that ensure consistency with the underlying physics of each system.
Visualization
A particularly useful tool for assessing the learned representation is to “traverse” the latent space and the space of reconstructions of the VAE [38, 32]. This can be achieved by generating synthetic data while interpolating over a specified generative factor, and setting the remaining generative factors to a constant reference value. The VAE is then evaluated on the generated data, yielding samples from the latent space, realizations of the reconstructed input $\hat{\bm{x}}$, and the mean physics-based and data-driven components $\hat{\bm{x}}_{\mathrm{p}}$ and $\hat{\bm{x}}_{\mathrm{d}}$. Each generative factor is linearly interpolated within the $1^{\mathrm{st}}$ and $99^{\mathrm{th}}$ percentiles of the corresponding ground truth distribution.
@@ 5.1 Beam case study
5.1.1 Case study description
The case study consists of a beam with fixed length $L=1.0$ m and a point load with magnitude $F=1.0$ N acting on an uncertain position $x_{\mathrm{F}}$ along the length of the beam. The material is linear elastic with uncertain Young’s modulus $E$, Poisson ratio $\nu=0.3$, area moment of inertia $I=2\cdot 10^{-6}$ m4 and cross-sectional area $A=2.4\cdot 10^{-3}$ m2. The rotational stiffness of the right-hand side support is temperature dependent, with the dependence modeled as an increase in the rotational stiffness of the support at lower temperatures. The relationship between the temperature and the support rotational stiffness is formulated as $\log k_{\mathrm{r}}=8-\frac{10}{1+e^{-T/2}}$. The beam is subject to variability in the vertical stiffness of the right-hand side support, e.g. due to damage or a deficiency of the support, simulated as a translational spring boundary condition with stiffness $k_{\mathrm{v}}$. This quantity can span several orders of magnitude, and therefore we parametrize the model using $\log k_{\mathrm{v}}$ instead. The beam is equipped with $d_{\mathrm{x}}=32$ sensors measuring the vertical displacement, equally spaced along the length as shown in Figure 1(a).
The Young’s modulus $E$ and the position of the point load $x_{\mathrm{F}}$ are considered as uncertain latent variables, such that $\bm{z}_{\mathrm{x}}=(E,x_{\mathrm{F}})$. It is assumed that the temperature is an observed domain variable such that $\bm{c}=(T)$, and that the vertical spring log-stiffness $\log k_{\mathrm{v}}$ is taken as a class variable representing damage in the structure, such that $\bm{y}=(\log k_{\mathrm{v}})$. Since the class variable $\bm{y}$ represents damage in the structure it will not be quantitatively measurable. In a realistic scenario, observations of the condition of the support on a qualitative scale (e.g. from 0 representing no damage to $5$ denoting a fully damaged support) might be available. In this example we simplistically consider $\bm{y}$ as the ground truth value of $\log k_{\mathrm{v}}$ with some added noise. The variable symbols, units, types, as well as the prior distributions over the physics-grounded latent variables and the ground truth distributions of the generative factors used to generate the training data are summarized in Table 1. We additionally provide a reference value which is used to produce the figures as discussed in Section 5. To ensure physical consistency and to avoid numerical issues, the Young’s modulus is truncated below a small positive value, and the load position $x_{\mathrm{F}}$ is restricted to the range $(0,1)$.
@@ 5.1.1 Case study description
where $b=L-x_{\mathrm{F}}$, and the non-bold $x$ refers to the position along the beam. This nominal model represents the beam in the undamaged condition at a reference temperature, and is directly incorporated in the physics-based branch of the VAE decoder.
@@
Following the procedure described in Section 5, the full model (trained on input-output pairs from an FE-based simulator) is used to produce synthetic data by first drawing samples of the input parameters from the ground truth distribution, and subsequently contaminating the resulting model predictions with zero-mean Gaussian white noise with standard deviation $\sigma_{\mathrm{x}}=0.02$ m. The dataset used to train the VAE is composed of $N_{\mathrm{train}}$ measurements of the beam displacement $\bm{x}=\{\bm{x}_{i}\}_{i=1}^{N_{\mathrm{train}}}$ where each element $\bm{x}_{i}$ is a vector of length $d_{\mathrm{x}}=32$. The domain and class observables $\bm{c}=\{\bm{c}_{i}\}_{i=1}^{N_{\mathrm{train}}}$ and $\bm{y}=\{\bm{y}_{i}\}_{i=1}^{N_{\mathrm{train}}}$ are obtained as the ground truth values used to generate the dataset, with the addition of i.i.d. samples of Gaussian white noise with standard deviations of $\sigma_{c}=\sigma_{y}=0.02$ respectively. The dimensionality of the domain and class latent variables is taken as $d_{z_{\mathrm{c}}}=d_{z_{\mathrm{y}}}=2$.
5.1.2 Qualitative assessment of disentanglement
@@
After training, the disentanglement between physics-grounded and data-driven components is qualitatively assessed by examining the latent space and samples of the reconstructed response of the beam. The predicted physics-based $\hat{\bm{x}}_{\mathrm{p}}$ and data-driven $\hat{\bm{x}}_{\mathrm{d}}$ components, as well as the combined prediction $\hat{\bm{x}}$ are shown in Figure 5. It can be observed that the data-driven component of the reconstruction $\hat{\bm{x}}_{\mathrm{d}}$ (middle row) is invariant to changes in $E$ and $x_{F}$ contributing only a constant deformed shape to the total predicted response. On the other hand, the physics-based model captures variability in both the physics-grounded generative factors $\bm{s}_{\mathrm{x}}=(E,x_{\mathrm{F}})$, but also domain and class generative factors $\bm{s}_{\mathrm{c}}=(T)$ and $\bm{s}_{\mathrm{y}}=(\log k_{\mathrm{v}})$. In contrast to the behavior of the unconstrained VAE, presented in Section 2.4, here the model preferentially utilizes the known physics. Only the variability in the measured response due to $\log{k_{\mathrm{v}}}$ and $T$ that can not be captured by the physics-based model is accounted for by the data-driven part of the decoder, indicating that the model can disentangle components of the response that can be attributed to the known physics from those that cannot. A key aspect of the adversarial training is the degree to which it allows interaction between the physics-based and data-driven components of the prediction. In this case study, the additional displacement of the beam due to the reduced vertical stiffness of the right-hand side support will also depend on the load position $x_{\mathrm{F}}$. More positive values of $\lambda$ tend to prevent the model from capturing this interaction, whereas more negative values enable it but may result in the data-driven components overriding the known physics.
@@
Figure 5: Mean prediction and $\pm 2\sigma$ uncertainty bounds for the physics-based $\hat{\bm{x}}_{\mathrm{p}}$ and data-driven $\hat{\bm{x}}_{\mathrm{d}}$ components, and combined prediction $\hat{\bm{x}}$ while traversing the generative factors. The input response measurements are denoted as dots in the bottom row.
To further highlight the impact of the GRL, the latent space traversals of the unconstrained model and the model trained adversarially are compared in Figure 6. Without adversarial training, the domain latent variables $\bm{z}_{\mathrm{c}}$ encode the variability in the load position $x_{\mathrm{F}}$ as shown in Figure 6(a), providing the data-driven decoder components with the information needed to reconstruct this component of the measured response and resulting in an entangled representation, as discussed in Section 2.4. In contrast, when $\lambda=1/256$ the adversarial training results in a posterior distribution over $\bm{z}_{\mathrm{c}}$ that is invariant to changes in $x_{\mathrm{F}}$. Instead, the variability is captured by the corresponding physics-grounded latent variable, as shown in Figure 6(b), indicating disentanglement of the physics-grounded and domain generative factors. The results shown previously suggest that the latent bottleneck architecture and GRL regularization result in a sparse and parsimonious representation of the physical system, and can yield domain and class latent variables that are invariant to changes in the underlying physics. The influence of the GRL hyperparameter $\lambda$ is further investigated in the oscillator case study presented below.
@@ 5.2 Oscillator case study
5.2.1 Case study description
This example demonstrates how the adversarial training prevents the model from compensating for all discrepancies between the physics-based model predictions and measurements. Suppose that a mass-spring-dashpot system undergoes damped harmonic motion, starting from an initial displaced position $x_{0}$, with no external excitation. It is assumed that each experiment is performed under varying temperature $T$, which is taken as the domain variable. The temperature affects the spring stiffness through the relationship $k(T)=k_{\mathrm{ref}}+\alpha_{\mathrm{T}}(T_{\mathrm{ref}}-T)$, with $T_{\mathrm{ref}}=20.0$ Co and $\alpha_{\mathrm{T}}=0.01$. The reference spring stiffness at $T_{\mathrm{ref}}=20.0$ Co is assumed known and equal to $k_{\mathrm{ref}}=1.0$ N/m. The mass $m$ is considered unknown and treated as a physics-grounded latent variable to be inferred from data. The viscous damping coefficient $c_{\mathrm{d}}$ is taken to define the class of the system. Finally, it is assumed that the observations are subject to an unknown confounding influence in the form of small random perturbations of the initial displacement $x_{0}$. A summary of the generative factors, the prior distribution and the ground truth distribution is provided in Table 2.
@@ 5.2.1 Case study description
Each triplet of observations is composed of a noisy displacement time series, and noisy measurements of the viscous damping coefficient $c_{\mathrm{d}}$ and temperature $T$, which are considered as class and domain variables respectively such that $\bm{c}=(T)$ and $\bm{y}=(c_{\mathrm{d}})$. Training and validation datasets are generated by drawing samples from the ground truth distribution and generating the oscillator displacement time-series using the full model, trained on input-output pairs simulated using the equation of motion shown in Equation 5.2. Each of the measured time-series is a vector of $64$ measurements, equally spaced within a time interval $t\in[0,10]$ s. The synthetic response measurements are subsequently contaminated with i.i.d. realizations of Gaussian white noise with standard deviation $\sigma_{\mathrm{x}}=0.01$ m. The standard deviation of the measurement uncertainty of the domain and class observables are taken as $\sigma_{\mathrm{c}}=0.01$ and $\sigma_{\mathrm{y}}=0.01$ respectively. To ensure that the model has sufficient capacity to learn the unknown confounding influence if the adversarial training were not present, the dimensionality of the latent space is specified to be significantly larger than the number of ground truth generative factors. The domain and class latent space dimensions are taken as $d_{z_{\mathrm{c}}}=d_{z_{\mathrm{y}}}=4$.
5.2.2 Model behavior in the presence of unknown confounders
The proposed model with no adversarial training ($\lambda=-1$) is trained and evaluated on the synthetic example. The reconstructed response obtained from a traversal of the initial displacement $x_{0}$, shown in 7(a), highlights another issue that occurs when combining physics-based and data-driven components in VAE: Although the variability in $x_{0}$ can not be accounted for by the physics-based model, and there is no information in the domain or class variables regarding the value of $x_{0}$, the lack of regularization results in a model that is free to capture the components of the measured displacement stemming from the variability in the initial displacement $x_{0}$. Although in this case the effect is benign, in more complex physical systems it can result in the model learning unknown confounding influences in a non-interpretable black-box manner. When the GRL regularization is utilized (7(b)), the data-driven encoder is unable to capture the variability in $x_{0}$, depriving the data-driven decoder from the information needed to reconstruct this component of the input measurements.
(a) Mean and $\pm 2\sigma$ uncertainty bounds of the reconstructed response without adversarial training ($\lambda=-1$).
(b) Mean and $\pm 2\sigma$ uncertainty bounds of the reconstructed response with adversarial training ($\lambda=1/128$).